OpenStudy (anonymous):

Simplify the expression: 4^2 + 8 / 7

OpenStudy (michele_laino):

here are some steps: \[\Large {4^2} + \frac{8}{7} = 16 + \frac{8}{7} = \frac{{\left( {16 \cdot 7} \right) + 8}}{7} = ...?\] please, for last step, keep in mind that the common denominator is 7